# Compactifications of phylogenetic systems and electrical networks

### v1

@@ 1. Introduction
Spaces of circular split systems are better known via phylogenetics, as in the book by Steele [36], and in recent papers by Catanzaro, Gambette, Balvočiūtė, Bryant, and Spillner [6, 22, 2]. Devadoss and Petti [11] showed that the space of circular split systems is an intuitive extension of the Billera-Holmes-Vogtmann space of phylogenetic trees [3], exhibiting a natural projection into the compactification of the real moduli space of curves [8, 9]. It is studied as a simplicial complex by Terhorst [37], and is related to the BME polytope and STSP polytope by Devadoss, Forcey, Scalzi, and Durell [10, 13, 16].
@@
Finally, Section 5 shows enumeration results (summarized by Table 1, page 1) across the spectrum of spaces and maps between them (showcased by Figure 6, page 6). Theorem 1 has an immediate application for counting the faithful split systems, resulting in the following (summarizing Corollary 22, and proven in Section 5):
@@
where $Pt_{n}$ is the set of Ptolemy diagrams.
@@ 5.2. Corollaries
Moreover, we can also count the compactified electrical networks, albeit more slowly than with the double factorial, using the same arguments as in the proof of Theorem 17:
@@ 5.3. Plabic Tilings and Ranges
The total number of (noncompact) unweighted split systems in the image $\mathbf{\xi}(\mathbf{\Omega}_{n})$ in $\mathbf{\Psi}_{n}$, which form a subcomplex of $\mathbf{\Psi}_{n},$ can be counted using a formula found in [26]. There, the Ptolemy diagrams are described as polygons with a subset of diagonals, obeying the rule that for every pair of crossing diagonals in a diagram, all the other diagonals using their four endpoints (edges of a quadrilateral) must be included in that diagram. [^12] The Ptolemy diagrams correspond to torsion pairs in the cluster category of type $A_{n}$ [26]. This rule precisely describes the Kron reductions of circular planar electrical networks (in clockwise order), with the extra requirement that the boundary edges (between consecutive nodes on the boundary) all be included. The Ptolemy diagrams $Pt_{n}$ of the $(k+3)$-gon are counted in [26] by
@@
resulting in the OEIS sequence [A181517]. Note that the plabic tilings of a single polygon with $n$ sides are in direct bijection with the Ptolemy diagrams of size $n$.
We wish to count the cells in the image $\mathbf{\xi}(\mathbf{\Omega}_{n}).$ The only difference between the polygonal pictures of split systems in the image of $\mathbf{\xi}$ and the Ptolemy diagrams is that the boundary edges (trivial splits) of the former can be included or excluded, when they are not part of a clique of size 4 or more. Denoting the number of these optional trivial splits for a given diagram $s$ by $t_{s}$ we have:
$|\mathbf{\xi}(\mathbf{\Omega}_{n})|\ =\ \sum_{s\in Pt_{n}}2^{t_{s}}.$
@@
When $n=4$, the four Ptolemy diagrams are $Pt_{4}=$[Uncaptioned image]. Thus we have $|\mathbf{\xi}(\mathbf{\Omega}_{4})|=2^{0}+2^{4}+2^{4}+2^{4}=49,$ as seen in Figure 3, inside the dashed line. For $n=1,\dots,5$ we have $|\mathbf{\xi}(\mathbf{\Omega}_{n})|=1,2,8,49,373$, respectively.
@@
For instance, when $n=4,$ $(5(49)+10(8)2+10(2)3+10(4)+5(1))/5=102.$ A summary of these results is given by Table 1.

### v2

@@ 1. Introduction
Spaces of circular split systems are better known via phylogenetics, as in the book by Steel [36], and in recent papers by Catanzaro, Gambette, Balvočiūtė, Bryant, and Spillner [6, 22, 2]. Devadoss and Petti [11] showed that the space of circular split systems is an intuitive extension of the Billera-Holmes-Vogtmann space of phylogenetic trees [3], exhibiting a natural projection into the compactification of the real moduli space of curves [8, 9]. It is studied as a simplicial complex by Terhorst [37], and is related to the BME polytope and STSP polytope by Devadoss, Forcey, Scalzi, and Durell [10, 13, 16].
@@
Finally, Section 5 proves enumeration results (summarized by Table 1, page 1) across the spectrum of spaces and maps between them (showcased by Figure 6, page 6). Theorem 1 has an immediate application for counting the faithful split systems, resulting in the following (expanding Corollary 22) :
@@
where $j_{0},\dots,j_{n}$ are ordered non-negative integers, $\mathcal{P}_{j_{i}}$ is the set of Ptolemy diagrams (from [26]) with $j_{i}$ vertices, and $t(s)$ is the number of optional edges of $s$.
@@ 5.2. Corollaries
Moreover, we can also count the compactified electrical networks, using the same arguments as in the proof of Theorem 17, or via a comment of P. Hanna in entry [A111088] of [oeis]:
@@ 5.3. Plabic Tilings and Ranges
The total number of (noncompact) unweighted split systems in the image $\mathbf{\xi}(\mathbf{\Omega}_{n})$ in $\mathbf{\Psi}_{n}$, which form a subcomplex of $\mathbf{\Psi}_{n},$ can be counted using a formula found in [26]. There, the Ptolemy diagrams are described as polygons with a subset of diagonals, obeying the rule that for every pair of crossing diagonals in a diagram, all the other diagonals using their four endpoints (edges of a quadrilateral) must be included in that diagram. [^12] The Ptolemy diagrams correspond to torsion pairs in the cluster category of type $A_{n}$ [26]. This rule precisely describes the Kron reductions of circular planar electrical networks (in clockwise order), with the extra requirement that the boundary edges (between consecutive nodes on the boundary) all be included. The Ptolemy diagrams $\mathcal{P}_{n}$ of the $n$-gon for $n=k+3$ are counted in [26] by
@@
resulting in the OEIS sequence [A181517]. Note that the plabic tilings of a single polygon with $n$ sides are in direct bijection with the Ptolemy diagrams of size $n$, for $n\geq 3$. We also include the unique diagrams of size $n=2$, 1, and 0: a single edge (a trivial split), a single vertex, and the empty diagram.
We wish to count the cells in the image $\mathbf{\xi}(\mathbf{\Omega}_{n}).$ The only difference between the polygonal pictures of split systems in the image of $\mathbf{\xi}$ and the Ptolemy diagrams is that the boundary edges (trivial splits) of the former can be included or excluded, when they are not part of a clique of size 4 or more. Denoting the number of these optional trivial splits for a given diagram $s$ by $t(s)$ we have:
$|\mathbf{\xi}(\mathbf{\Omega}_{n})|\ =\ \sum_{s\in\mathcal{P}_{n}}2^{t(s)}.$
@@
When $n=4$, the four Ptolemy diagrams are $Pt_{4}=$[Uncaptioned image]. Thus we have $|\mathbf{\xi}(\mathbf{\Omega}_{4})|=2^{0}+2^{4}+2^{4}+2^{4}=49,$ as seen in Figure 3, inside the dashed line. For $n=1,\dots,6$ we have $|\mathbf{\xi}(\mathbf{\Omega}_{n})|=1,2,8,49,373,3196$ respectively.
@@
Expanding this formula gives Corollary 2. For example, when $n=4,$ we have the total:
$(1/5)(5(49)+20(8)+10(4)+30(2)+5(1))=102.$ A summary of early results is given by Table 1.